Determine whether the system is consistent, inconsistent, or dependent.
3x - 3y = -9
9x - 9y = -27

Found 2 solutions by stanbon, hokies:
Answer by stanbon(75887)

Determine whether the system is consistent, inconsistent, or dependent.
3x - 3y = -9
9x - 9y = -27
-------------
Multiply thru the 1st eqution by 3 to get:
9x-9y = -27
---
Subtract that from the 2nd equation and you get:
0 = 0
---
That means the system is dependent (one of the equations
is a multiply of the other).
====================================
Cheers,
Stan H.

Answer by hokies(65)

dependent. eqn 2 is 3 times eqn 1.